# Jialiang Fan Franklin College of Arts and Sciences

### v1

@@ 3.3 Comparative experiments
We conducted comparative experiments using classical computer vision models, including VGG-16 [9] and ResNet18 [10], on our augmented dataset. The results, presented in Table 3, indicate that both models perform exceptionally well on the testing dataset. Specifically, ResNet18 achieves an accuracy of 0.999 on Classifier 1, while VGG16 reaches perfect accuracy, correctly classifying all test images on Classifier 1. The performance on Classifier 2, though slightly lower, remains impressive: ResNet18 and VGG16 achieve accuracies of 0.997 and 0.995, respectively. Overall, both ResNet18 and VGG16 slightly outperform the Vision Transformer (ViT) model.
Table 3: Performance comparison of two classifiers with and without data augmentation
Classifier Encoder Accuracy Precision Recall F1 Score Specificity
Classifier 1 ViT 0.989 0.997 0.979 0.988 0.998
Classifier 1 ResNet18 0.999 1 0.998 0.999 1
Classifier 1 VGG16 1 1 1 1 1
Classifier 2 ViT 0.971 0.97 0.967 0.968 0.988
Classifier 2 ResNet18 0.997 1 0.993 0.996 1
Classifier 2 VGG16 0.995 1 0.988 0.994 1
@@ 4 Discussion
Based on the built dataset, we propose the dual-task vision transformer (DTViT) model based on the vision transformer for dual-task classification. The proposed model is composed of an encoder to extract information from CT images and two decoders for different classification tasks, i.e., classification of normal and ICH images and classification of three types of ICH based on the location of the hematoma. The proposed DTViT achieves 99.7% of the training accuracy and 99.5% of the testing accuracy on the augmented dataset. We also compared the proposed model with several state-of-the-art models, i.e., Resnet18, Resnet34, Vgg16, and Alexnet, and the results show that the proposed model achieves a desirable result.
@@ 5 Conclusion
In this paper, we have constructed a dataset for Intracerebral Hemorrhage (ICH) based on real-world clinical data. The dataset comprises CT images that have been initially processed and manually classified by medical experts into categories of normal and ICH images. Moreover, the ICH images have been categorized into three types-Deep, Subcortical, and Lobar-according to the hemorrhage’s location.
Additionally, we have introduced a dual-task vision Transformer (DTViT) aimed at the classification of Intracerebral hemorrhage. This innovative neural network model incorporates an encoder, which utilizes the cutting-edge vision Transformer architecture, and two decoders that are designed to classify images as ICH or normal and to determine the hemorrhage type. Our experiments have demonstrated that the DTViT has achieved a remarkable accuracy rate of 99.7% on the test data. We have also compared the DTViT with traditional models such as ResNet, VGG, and AlexNet.

### v2

@@ 3.4 Comparative experiments
To better demonstrate the performance of the DTViT, we conducted comparative experiments using classical CNN models, including ResNet18[22], AlexNet[23], SqueezeNet[24], and DenseNet[25] on our augmented dataset. The comparative results are shown in Table 4. Specifically, the proposed model, i.e. DTViT, achieves the lowest loss of $0.0102$ and the highest accuracy of $0.9988$ over these models. In addition, the DTViT also ranks first on other evaluation metrics, except its precision is slightly lower than DenseNet’s.
Table 4: Overall performance comparisons with existing models.
Model Loss Accuracy Precision Recall F1 Score Specificity
Proposed 0.0102 0.9988 0.9991 0.9983 0.9987 1
ResNet18[22] 0.0170 0.9984 0.9989 0.9983 0.9986 0.9993
AlexNet[23] 0.0142 0.9976 0.9989 0.9983 0.9986 0.9994
SqueezeNet[24] 0.1244 0.9775 0.990 0.9889 0.9895 0.9993
DenseNet[25] 0.0119 0.9976 1 0.9964 0.9981 1
@@ 4 Discussion
Based on the built dataset, we propose the dual-task vision transformer (DTViT) model based on the vision transformer for dual-task classification. The proposed model is composed of an encoder to extract information from CT images and two decoders for different classification tasks, i.e., classification of normal and ICH images and classification of three types of ICH based on the location of the hematoma. The proposed DTViT achieves 99.7% of the training accuracy and 99.88% of the testing accuracy on the augmented dataset. We also compared the proposed model with several state-of-the-art models, i.e., Resnet18, SqueezeNet, and Alexnet, and the results show that our model achieves the best performance over these models.
@@ 5 Conclusion
In this paper, we have constructed a dataset for Intracerebral Hemorrhage (ICH) based on real-world clinical data. The dataset comprises CT images that have been initially processed and manually classified by medical experts into categories of normal and ICH images. Moreover, the ICH images have been categorized into three types, Subcortical, and Lobar, according to the hemorrhage’s location.
Additionally, we have introduced a dual-task vision Transformer (DTViT) aimed at the classification of Intracerebral hemorrhage. This innovative neural network model incorporates an encoder, which utilizes the cutting-edge vision Transformer architecture, and two decoders that are designed to classify images as ICH or normal and to determine the hemorrhage type. Our experiments have demonstrated that the DTViT has achieved a remarkable accuracy rate of $99.88\%$ on the test data. We have also compared the DTViT with traditional models such as Resnet18, SqueezeNet, and Alexnet.